# Functional bottlenecks can emerge from non-epistatic underlying traits

### v1

@@ II.4 Results for the calibrated model
[…]
@@
In the following, we thus choose $L=500$ and the parameters corresponding to the black dots in Fig. 3 as the final parameters for which we report the results: Gaussian $P(h)$ with $p=0.26$ and $E_{T}=2.0$ for the tuning procedure, and Pareto cutoff $P(h)$ with $p=0.25$ and $E_{T}=1.1$. Keeping these two models fixed, we generated many ($\sim 20000$) random instances of the model and studied the distribution of the relevant quantities that characterize the topology of the fitness landscape. An example of a ‘good’ topology, closely resembling the one found experimentally, has already been given in Fig. 2. We want to quantify the probability of generating such a topology in the toy model.
We first focus on the properties of the two wild types, shown in Fig. 4 for both models. More specifically, Figs. 4a,e show the histograms of $M$, the total number of mutations separating the two wild types. We note that, by construction, $M\geq 2$, and we fixed $\langle M\rangle=8$. We see that the distribution is quite broad, leading to a significant fraction of instances with $M$ as large as 20, and resembles that shown for elastic networks in Ref. Alqatari and Nagel (2024). However, when we investigate the mutational paths, we restrict the analysis to instances with $M\leq 14$ for computational tractability, as in Ref. Alqatari and Nagel (2024). Figs. 4b,f show the histograms of $E^{WT}_{B}/E_{T}$ for the blue wild type (similar results are obtained for the red wild type due to the inherent symmetry of the model). We recall that, by construction of the training procedure, $E^{WT}_{B}/E_{T}\geq 1$. We observe that the histogram is peaked slightly above the minimal value, which confirms that the tuning procedure produces $E^{WT}_{B}\sim E_{T}$ and $E^{WT}_{R}\sim-E_{T}$. Figs. 4c,g show the histograms of the number of greedy steps in the tuning procedure, which is peaked around one, as expected. Finally, Figs. 4d,h show the histograms of the $M$ SMEs, here called $\widetilde{h}_{i}$, that are selected by the tuning procedure, hence characterize a posteriori (after tuning) the space of intermediates between the wild type. We observe in both cases a strongly bimodal distribution, which reflects the choice of the tuning procedure: mutations selected by the random steps tend to be small, while mutations selected by the greedy steps tend to be much larger (in a way that can be quantified by extreme value statistics, see the SI).
Next, we analyze the statistical properties of the space of paths connecting the wild types (Fig. 5). We recall that in the calibrated models we find $E_{C}<E^{WT}_{\rm min}$ with probability one. The space of paths thus contains a single genotype through which all paths must pass: raising $E_{C}$ would remove this genotype from the set, and disconnect the wild types. A bottleneck is thus observed. We note that the jumper genotype, in most cases, is connected by a single mutation to only one genotype carrying the other phenotype, as in Fig. 2. However, in a few cases, we also observe situations where the jumper genotype is connected to a few genotypes carrying the other phenotype. In Fig. 5a,d we show the histograms of $E_{C}/E^{WT}_{\rm max}$, separately for instances with given $M$. We observe that the histogram is almost independent of $M$ (with more variability observed for the Pareto cutoff model), and peaked around $E_{C}/E^{WT}_{\rm max}\sim 1/2$ as discussed above. Fig. 5b,e shows the histogram of the location of the jumper, called $j$, along the mutational path. We observe that the jumper is often located mid-way between the wild types, leading to a peaked histogram around $j/M\sim 1/2$. The peak (slowly) becomes sharper upon increasing $M$. The qualitative shape of the possible topologies produced by the model is indicated in Fig. 5g. Finally, we can compute the number of paths that remain viable at $E_{C}$. The average of the logarithm of this number is shown in Fig. 5c,f as a function of $M$. We find that it grows almost linearly at large $M$, indicating that the number of paths grows exponentially in $M$. This suggests that there are exponentially many ways to reach the bottleneck, which could thus maintain a good evolutionary accessibility from both wild types Alqatari and Nagel (2024).
@@
Figure 5: Statistical properties of the space of paths for two different calibrated model, the Gaussian model (a-c) and the Pareto cutoff model (d-f), with the same parameters as in Fig. 4. (a), (d) Distribution of values of $E_{C}/E^{WT}_{\rm max}$, separately for several values of $M$. The mean and the median (dashed and dot-dashed vertical lines, respectively) are computed with respect to the full distribution (with all $M$ values). (b), (e) Probability distribution of the jumper position $j$ along the path, for different values of $M$. (c), (f) Average of logarithm of the number of paths that remain viable at $E_{C}$, as a function of $M$. (g) Some examples of topologies, each associated to a schematic representation used in panel b.
@@
We conclude that our toy model, which features a global non-linearity on top of an underlying additive phenotype, with properly calibrated parameters, generates topologies that are very close to those observed experimentally. An example is shown in Fig. 2, to be compared with the result obtained re-analyzing the experimental data from Ref. Poelwijk et al. (2019), shown in Fig. 1d. The model generates an ensemble of random fitness functions Pahujani and Krug (2025), whose topology is likely (with probability one) characterized by a bottleneck, i.e., by a genotype through which all paths connecting the two wild types must pass. Of course, the bottleneck can be made less pronounced if one allows for less stringent selection, i.e., reduces $E_{C}$ Alqatari and Nagel (2024). The bottleneck is more likely found mid-way along the path. These features are essentially independent of $M$, i.e., the number of mutations separating the two wild types, at least in the limited range investigated here.
@@
We note that there is quite a lot of flexibility in the choice of parameters. The initial size $L$ can be varied in a broad interval without affecting much the results (see SI). The tuning parameter $E_{T}$ has been chosen here in such a way to obtain the desired average of $M$, but other choices are possible. The value of $p$, i.e., the probability of making a greedy step when constructing the wild types, can be varied in a relatively wide range, provided it is neither too small (otherwise $E_{T}$ is too small and the wild types end up overshooting the fitness threshold) nor too large (otherwise $\langle E_{C}/E^{WT}_{\rm max}\rangle$ decreases. Finally, the a priori distribution $P(h)$ that characterizes the input distribution of SMEs can be varied essentially at will in a broad range (see also the SI for other choices of the input distribution). The tuning procedure used to construct the wild types ensures that the SMEs selected a posteriori are broadly distributed (Fig. 4d,h), which is a necessary condition for the emergence of a bottleneck. We thus believe that our results are robust and do not depend crucially on the specific choices made during calibration.
@@ III Discussion
In this work, we have demonstrated that functional bottlenecks in protein fitness landscapes can emerge even in the absence of network epistasis, solely due to a non-linear mapping between an underlying additive phenotype and fitness. By constructing and analyzing a simple model of global epistasis, we have shown that the emergence of bottlenecks is closely tied to the heterogeneity of the single-mutation effects selected during the construction of the wild types. When mutational effects are broadly distributed, evolutionary paths between distinct phenotypes are constrained, leading to the formation of narrow evolutionary corridors.
@@
Our findings provide a novel perspective on how fitness landscape topologies arise and suggest that strong constraints on evolutionary accessibility can exist even in systems governed by simple, global epistatic interactions De Visser and Krug (2014); Tenaillon (2014); Pahujani and Krug (2025); Hwang et al. (2017); Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Reddy and Desai (2021); Phillips et al. (2021); Johnson et al. (2023); Buda et al. (2023). This has important implications for protein evolution, as it suggests that functional bottlenecks may not necessarily indicate complex, higher-order interactions but could instead be a natural consequence of non-linear selection on an underlying additive phenotype. Moreover, our model successfully captures key features observed in experimental data, supporting the idea that broad mutational effect distributions play a central role in shaping evolutionary trajectories.
@@
The model is simple enough that it could possibly be solved analytically, thus eliminating the limitation on the number of mutations $M$ separating the two wild types, using techniques from probability and statistical physics De Visser and Krug (2014); Hwang et al. (2017); Pahujani and Krug (2025). Future work could explore this possibility to extend our approach to larger mutational spaces, incorporate additional biological constraints, and investigate the robustness of these results across different fitness functions and evolutionary pressures, possibly with more than two phenotypes.
@@ IV.1 Genotype representation and underlying additive phenotype
The first step consists in choosing a value of $L$ and considering a genotype ${\bf a}=(a_{1},a_{2},\dots,a_{L})$ as a binary sequence of length $L$, where each site $i$ can take values $a_{i}\in\{0,1\}$. The genotype-dependent underlying additive phenotype is given by:
@@ IV.2 Fitness function and phenotypic classes
The global epistatic fitness function is a non-linear transformation of $E(a)$, defined separately for the blue (B) and red (R) phenotypes:
@@
IV.3 Wild type construction
The red and blue wild types, ${\bf a}_{R}$ and ${\bf a}_{B}$, are independently generated starting from the ancestral genotype ${\bf a}=(0,0,\dots,0)$ by sequentially introducing mutations. At each step, with probability $p$, the mutation with the largest contribution to $E$ in the desired direction is chosen (greedy step), and with probability $1-p$, a random site is mutated (random step). The process is iterated until $E^{WT}_{R}=E({\bf a}_{R})<-E_{T}$ and $E^{WT}_{B}=E({\bf a}_{B})>E_{T}$.
@@ IV.4 Mutational paths and bottleneck characterization
The mutational distance $M$ between ${\bf a}_{R}$ and ${\bf a}_{B}$ is defined as the number of differing sites. All $M!$ possible evolutionary paths between the wild types are considered, and a value $E_{C}$ is determined such that at least one path of single mutations exists, maintaining $|E({\bf a})|>E_{C}$ at all intermediate steps. The distribution of $E_{C}$ and the likelihood of a single critical ‘jumper’ mutation define the presence and severity of the bottleneck.
@@ Acknowledgements.
We warmly thank Sid Nagel and Martin Weigt for providing very useful feedback and inspiration all along the development of this work. This research has been supported by first FIS (Italian Science Fund) 2021 funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from MUR, Italian Ministry of University and Research and from the PRIN funding scheme (2022LMHTET - Complexity, disorder and fluctuations: spin glass physics and beyond) from MUR, Italian Ministry of University and Research.
SUPPLEMENTARY INFORMATION
@@
The experimental measurements that were analyzed in the main text were obtained from the public repository associated with Ref. Poelwijk et al. (2019). These are raw sequencing data capturing the genotype variants present in the red and blue channel after two-color Fluoresce Activated Cell Sorting. Following the procedure detailed in Ref. Poelwijk et al. (2019) (see the original paper for details), sequences with low quality scores and noisy measurements were filtered out. The enrichment scores for each genotype a were computed counting the occurrence of each variant in the two channels normalized to its abundance in the input channel, i.e. before sorting. We further normalized the enrichment scores with respect to the wild types, such that $F_{R}(\textbf{a}_{R})=F_{B}(\textbf{a}_{B})=1$, and afterwards removed global epistasis via a non-linear function
@@
as proposed by the authors of Ref. Poelwijk et al. (2019), separately for the blue (${f}=B$) and red (${f}=R$) fluorescence.
@@
Figure S1: (a) Scatter plot of the enrichment scores in the blue and the red channel, $\epsilon_{B}$ and $\epsilon_{R}$, not normalized with respect to the wildtypes. For the blue wildtype $\epsilon_{B}=1.58$, $\epsilon_{R}=0.09$. For the red wildtype $\epsilon_{B}=0.17$, $\epsilon_{R}=0.73$. (b) Histogram of the values of $\widetilde{E}$ obtained for each of the $2^{13}$ variants. The red and blue lines correspond respectively to the wild type values $\widetilde{E}^{WT}_{R}$ and $\widetilde{E}^{WT}_{B}$. (c) Topology of the space of paths obtained keeping only genotypes $\mathbf{a}$ with $|\widetilde{E}(\mathbf{a})|>\widetilde{E}_{C}$, where $\widetilde{E}_{C}=0.56$ (dashed lines) is the largest possible value such that the red and blue wild types remain connected. The lower panel reports the values of $\widetilde{E}(\mathbf{a})$ for each functional genotype (blue dots for $\widetilde{E}>\widetilde{E}_{C}$ and red dots for $\widetilde{E}<-\widetilde{E}_{C}$) as a function of the number of mutations $m$ from the blue wild type, with the gray lines connecting pairs of genotypes that differ by a single mutation. The upper panel shows the resulting graph of connections.
@@
In the main text, and as discussed in Sec. S1.1, we normalized the enrichment scores to the wild type values, a choice that was not made in the original Ref. Poelwijk et al. (2019). Using normalized enrichment scores, we find that no path of single mutations exists that connects the two wild types while always maintaining $|E(\textbf{a})|\geq E^{WT}$, hence we need to consider $E_{\rm{th}}<E^{WT}$ in order to preserve the connection between the wild types, as discussed in the main text. Instead, Ref. Poelwijk et al. (2019) shows a topology of the space of paths connecting genotypes with $|E(\textbf{a})|\geq E_{\rm{th}}$ and $E_{\rm{th}}=E^{WT}_{\rm{min}}$. This discrepancy arises entirely from the choice of not normalizing the phenotypic trait.
The space of paths presented in Ref. Poelwijk et al. (2019) is constructed by defining a phenotype $E=\sqrt{\epsilon^{2}_{B}+\epsilon^{2}_{R}}$, where $\epsilon_{B}$ and $\epsilon_{R}$ are the enrichment scores in the two channels, which are not normalized with respect to the enrichment scores of the red and blue wild types. However, they are rescaled to reflect the experimental brightness ratio between the parental blue and red genotypes, which is $0.172$, and subsequently transformed using the non-linear function in Eq. S1 to remove global epistasis. To carry out a proper comparison with our analysis, we reproduce the space of paths of Ref. Poelwijk et al. (2019) modifying their definition of phenotype to $\widetilde{E}=\epsilon_{B}-\epsilon_{R}$. Fig. S1a shows that this is a minor change: because $\epsilon_{R}\sim 0$ when $\epsilon_{B}>0$ and vice versa, the two phenotypes are essentially exclusive and as a consequence $E=\sqrt{\epsilon^{2}_{B}+\epsilon^{2}_{R}}\sim\max(\epsilon_{B},\epsilon_{R})\sim|\widetilde{E}|$. Hence, $\widetilde{E}$ is numerically very close to the score used in the original Ref. Poelwijk et al. (2019), while keeping the negative sign for the red phenotype.
@@
Compared to the normalized case shown in the main text, the resulting distribution of phenotypes, shown in Fig. S1b, is not symmetric around zero: the blue wildtype has $|\widetilde{E}^{WT}_{R}|=0.56$ and the red one is more then twice as large, with $\widetilde{E}^{WT}_{B}=1.49$. This asymmetry arises entirely from the absence of normalization. When looking for the largest value of the functionality threshold $\widetilde{E}_{\rm{th}}$ such that at least one viable mutational path exists between the two wild types, we find that $\widetilde{E}_{C}$ saturates to $\widetilde{E}^{WT}_{\rm{min}}$ and does not produce a topology with a single jump. The topology is the same as the one shown in Ref. Poelwijk et al. (2019), and we consistently find that $\widetilde{E}_{C}$ = $\widetilde{E}^{WT}_{\rm{min}}$ is the critical value above which any connection between the two wild types is lost. This is exactly what we would expect given the imbalance between the two wild types. However, given the imbalance between wild types, we believe that choosing a symmetric threshold, i.e. choosing the same $\widetilde{E}_{\rm{th}}$ for both functions, is not appropriate.
@@
We therefore believe that the choice of normalizing the enrichment scores to the wild types provides a better view of the global fitness landscape, with the distribution of phenotypic values centered around zero. This choice results in a critical threshold $\widetilde{E}_{C}<\widetilde{E}_{WT}$, as discussed in the main text.

### v2

@@ II.4 Results for the calibrated model
Figure 4: Distribution of some relevant quantities for two different calibrated models, the Gaussian model and the Pareto cutoff model. In the former, $L=500$ SMEs are extracted from a Gaussian distribution with unit variance and the tuning procedure has $p=0.26$, $E_{T}=2.0$; in the latter, $L=500$ SMEs are extracted from a distribution with Pareto tails decaying with $\alpha=0.7$ and a cutoff $|h|<2.0$, and the tuning procedure has $p=0.25$, $E_{T}=1.1$. In each plot the dashed and dot-dashed vertical lines represent the mean and the median of the data, respectively. (A) Distribution of the number of mutations. (B) Distribution of the value of the positive (blue) reference phenotype divided by $E_{T}$. (C) Distribution of the number of greedy steps needed to reach the target value $E_{T}$ when generating the two reference variants. (D) Distribution of SMEs $\tilde{h}_{i}$ selected by the tuning procedure to generate the two reference variants.
@@
In the following, we choose $L=500$ and the parameters corresponding to the black dots in Fig. 3 as the final calibrated parameters for which we report more detailed results: Gaussian $P(h)$ with $p=0.26$ and $E_{T}=2.0$ for the tuning procedure, and Pareto cutoff $P(h)$ with $p=0.25$ and $E_{T}=1.1$. Keeping these two models fixed, we generated many ($\sim 20000$) random instances of the model and studied the distribution of the relevant quantities that characterize the topology of the fitness landscape. An example of a ‘good’ topology, closely resembling the one found experimentally, has already been given in Fig. 2. We want to quantify the probability of generating such a topology in the stylized model.
@@
Figure 5: Statistical properties of the space of paths for two different calibrated model, the Gaussian model (A-C) and the Pareto cutoff model (D-F), with the same parameters as in Fig. 4. (A), (D) Distribution of values of $E_{C}/E^{\rm{ref}}_{\rm max}$, separately for several values of $M$. The mean and the median (dashed and dot-dashed vertical lines, respectively) are computed with respect to the full distribution (with all $M$ values). (B), (E) Probability distribution of the jumper position $j$ along the path, for different values of $M$. (C), (F) Average of logarithm of the number of paths that remain viable at $E_{C}$, as a function of $M$. (G) Some examples of topologies, each associated to a schematic representation used in panel b.
We first focus on the properties of the two reference variants, shown in Fig. 4 for both models. More specifically, Fig. 4a shows the histograms of $M$, the total number of mutations separating the two reference sequences. We note that, by construction, $M\geq 2$, and we fixed $\langle M\rangle=8$. We see that the distribution is quite broad, leading to a significant fraction of instances with $M$ as large as 20, and resembles that shown for elastic networks in Ref. Alqatari and Nagel (2026). However, when we investigate the mutational paths, we restrict the analysis to instances with $M\leq 14$ for computational tractability, as in Ref. Alqatari and Nagel (2026). Fig. 4b shows the histograms of $E^{\rm{ref}}_{B}/E_{T}$ for the blue reference variant (similar results are obtained for the red reference variant due to the inherent symmetry of the model). We recall that, by construction of the training procedure, $E^{\rm{ref}}_{B}/E_{T}\geq 1$. We observe that the histogram is peaked slightly above the minimal value, which confirms that the tuning procedure produces $E^{\rm{ref}}_{B}\sim E_{T}$ and $E^{\rm{ref}}_{R}\sim-E_{T}$. Fig. 4c shows the histograms of the number of greedy steps in the tuning procedure, which is peaked around one, as expected. Finally, Fig. 4d shows the histograms of the $M$ SMEs $\tilde{h}_{i}$ that are selected by the tuning procedure. These mutations characterize a posteriori (after tuning) the space of intermediates between the reference variants, and can be thought of as being those ‘fixed’ by evolution. We observe a strongly bimodal distribution P(h~)P($\tilde${h}) for both input distributions $P(h)$, which reflects the choice of the tuning procedure: mutations selected by the random steps tend to be small, while mutations selected by the greedy steps tend to be much larger (in a way that can be quantified by extreme value statistics, see the SI).
Next, we analyze the statistical properties of the space of paths connecting the reference variants (Fig. 5). We recall that in the calibrated models we find $E_{C}<E^{\rm{ref}}_{\rm min}$ with probability one. The space of paths thus contains a single genotype through which all paths must pass: raising $E_{C}$ would remove this genotype from the set, and disconnect the reference variants. A bottleneck is thus observed. We note that the jumper genotype, in most cases, is connected by a single mutation to only one genotype carrying the other phenotype, as in Fig. 2. However, in a few cases, we also observe situations where the jumper genotype is connected to a few genotypes carrying the other phenotype. In Fig. 5a,d we show the histograms of $E_{C}/E^{\rm{ref}}_{\rm max}$, separately for instances with given $M$. We observe that the histogram is almost independent of $M$ (with more variability observed for the Pareto cutoff model), and peaked around $E_{C}/E^{\rm{ref}}_{\rm max}\sim 1/2$ as discussed above. Fig. 5b,e shows the histogram of the location of the jumper, called $j$, along the mutational path. We observe that the jumper is often located mid-way between the reference variants, leading to a peaked histogram around $j/M\sim 1/2$. The peak (slowly) becomes sharper upon increasing $M$. The qualitative shape of the possible topologies produced by the model is indicated in Fig. 5g. Finally, we can compute the number of paths that remain viable at $E_{C}$. The average of the logarithm of this number is shown in Fig. 5c,f as a function of $M$. We find that it grows almost linearly at large $M$, indicating that the number of paths grows exponentially in $M$. This suggests that there are exponentially many ways to reach the bottleneck, which could thus maintain a good evolutionary accessibility from both reference sequences Alqatari and Nagel (2026).
@@
We conclude that our stylized model, which features a global non-linearity on top of an underlying additive phenotype, with properly calibrated parameters, generates topologies that are very close to those observed experimentally in Ref. Poelwijk et al. (2019) and Sec. II.1. An example is shown in Fig. 2, to be compared with the result obtained re-analyzing the experimental data from Ref. Poelwijk et al. (2019), shown in Fig. 1d. The model generates an ensemble of random fitness functions Pahujani and Krug (2025), whose topology is likely (with probability one) characterized by a bottleneck, i.e., by a genotype through which all paths connecting the two reference variants must pass. Of course, the bottleneck can be made less pronounced if one allows for less stringent selection, i.e., reduces $E_{C}$ Alqatari and Nagel (2026). The bottleneck is more likely found mid-way along the path. These features are essentially independent of $M$, i.e., the number of mutations separating the two reference variants, at least in the limited range investigated here.
@@
We note that there is quite a lot of flexibility in the choice of parameters. The initial size $L$ can be varied in a broad interval without affecting much the results (see SI). The tuning parameter $E_{T}$ has been chosen here in such a way to obtain the desired average of $M$, but other choices are possible. The value of $p$, i.e., the probability of making a greedy step when constructing the reference variants, can be varied in a relatively wide range, provided it is neither too small (otherwise $E_{T}$ is too small and the reference phenotypes end up overshooting the fitness threshold) nor too large (otherwise $\langle E_{C}/E^{\rm{ref}}_{\rm max}\rangle$ decreases). Finally, the a priori distribution $P(h)$ that characterizes the input distribution of SMEs can be varied essentially at will in a broad range (see also the SI for other choices of the input distribution). The tuning procedure used to construct the reference variants ensures that the SMEs selected a posteriori are broadly distributed, featuring the right balance between almost neutral SMEs and strongly beneficial or deleterious SMEs (Fig. 4d), which is a necessary condition for the emergence of a bottleneck. We thus believe that our results are robust and do not depend crucially on the specific choices made during calibration.
@@ III Discussion
In this work, we have demonstrated that functional bottlenecks in protein fitness landscapes can emerge even in the absence of network epistasis, solely due to a non-linear mapping between an underlying additive trait and fitness. To do so, we constructed and analyzed a simple stylized model of global epistasis, in which two reference variants are constructed from a common ancestor by selecting at random either a very beneficial mutation, or a neutral one, with the goal of acquiring two distinct new phenotypes. We have shown that the emergence of bottlenecks separating the two phenotypes is closely tied to the heterogeneity of the single-mutation effects selected during the construction of the reference variants. When mutational effects are heterogeneously enough, evolutionary paths between distinct phenotypes are constrained, leading to the formation of narrow evolutionary corridors.
@@
Our findings provide a novel perspective on how fitness landscape topologies arise and suggest that strong constraints on evolutionary accessibility can exist even in systems governed by simple, global epistatic interactions De Visser and Krug (2014); Tenaillon (2014); Pahujani and Krug (2025); Hwang et al. (2017); Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Reddy and Desai (2021); Phillips et al. (2021); Johnson et al. (2023); Buda et al. (2023). In fact, in real evolution experiments, one always has access to a single realization of the process. Our stylized model, following and simplifying the approach of Ref. Alqatari and Nagel (2026), constructs an ensemble of fitness landscapes and allows us to assess to what extent the realizations observed in a given experiment are ‘typical’ of a reasonable ensemble of possibilities. Moreover, despite being stylized, our model successfully captures key qualitative features observed in experimental data, supporting the idea that broad mutational effect distributions play a central role in shaping evolutionary trajectories.
Our results indicate that functional bottlenecks do not necessarily entail network epistasis. Such bottlenecks may instead arise because proteins, under strong evolutionary pressure to acquire new functions, tend to accumulate a few mutations that confer large benefits for the new function, alongside a handful of random neutral mutations introduced by the inherent stochasticity of evolution. Hence, functional bottlenecks may not necessarily indicate complex, higher-order interactions but could instead be a natural consequence of non-linear selection on an underlying additive trait.
Of course, network epistasis does exist in proteins, as it has been shown by a variety of approaches Lunzer et al. (2010); Olson et al. (2014); Figliuzzi et al. (2016); Rivoire et al. (2016); Poelwijk et al. (2016); Cocco et al. (2018); Starr et al. (2018); Russ et al. (2020); Ballal et al. (2020); Lyons et al. (2020); Miton et al. (2021); Bakerlee et al. (2022); Vigué et al. (2022); Chen et al. (2024). Our stylized model might serve as a null model to assess the relevance of network epistasis effects in the analysis of future experimental data.
We also stress that in different kinds of evolutionary dynamics, for instance neutral space evolution under weak selection Fantini et al. (2020); Stiffler et al. (2020); Bisardi et al. (2022), mutations are selected in a different way and our stylized model would likely not apply. The existence of functional bottlenecks during this kind of evolutionary dynamics is less established and might depend on subtle correlations between large groups of interacting mutations Rivoire et al. (2016); Bari et al. (2024); Rossi et al. (2025). A recent study of a combinatorial landscape separating naturally occurring bacterial transcription factor binding sites did not find a bottleneck Westmann et al. (2024).
@@
Our model is simple enough that it could possibly be solved analytically, using techniques from probability and statistical physics De Visser and Krug (2014); Hwang et al. (2017); Pahujani and Krug (2025). For instance, the tuning procedure consists in a random walk in the space of mutations, and preliminary results suggest that the probability of the resulting reference traits $E^{\rm{ref}}_{B}$ and $E^{\rm{ref}}_{R}$ can be written in simple form. Making some analytical progress would eliminate the limitation on the number of mutations $M$ separating the two reference variants. Future work could explore this possibility to extend our approach to larger mutational spaces, incorporate additional biological constraints, and investigate the robustness of these results across different fitness functions and evolutionary pressures, possibly with more than two phenotypes.
It should be noted that numerous techniques for extracting genotype-trait-fitness relationships directly from data – including modern machine learning frameworks – have emerged in what is now a rapidly expanding field (see e.g., Figliuzzi et al. (2018); Wang et al. (2018); Tubiana et al. (2019); Russ et al. (2020); Trinquier et al. (2021); Rives et al. (2021); Malbranke et al. (2023); Wagner (2024); Razo-Mejia et al. (2025)). The present work adopts a fundamentally different approach. Rather than attempting to fit specific empirical datasets, we propose here a simple stylized model, designed to identify relevant qualitative features and universal constraints in a generic way.
@@ IV.1 Genotype representation and underlying additive phenotype
The first step consists in choosing a value of $L$ and considering a genotype ${\bf a}=(a_{1},a_{2},\dots,a_{L})$ as a binary sequence of length $L$, where each site $i$ can take values $a_{i}\in\{0,1\}$. The genotype-dependent underlying additive trait is given by:
@@ IV.2 Fitness function and phenotypic classes
The global epistatic fitness function is a non-linear transformation of $E({\bf a})$, defined separately for the blue (B) and red (R) phenotypes:
@@
IV.3 Reference variants construction
The red and blue reference variants, ${\bf a}_{R}$ and ${\bf a}_{B}$, are independently generated starting from the ancestral genotype ${\bf a}=(0,0,\dots,0)$ by sequentially introducing mutations. At each step, with probability $p$, the mutation with the largest contribution to $E$ in the desired direction is chosen (greedy step), and with probability 1p1-p, a random site is mutated (random step). The process is iterated until $E^{\rm{ref}}_{R}=E({\bf a}_{R})<-E_{T}$ and $E^{\rm{ref}}_{B}=E({\bf a}_{B})>E_{T}$.
@@ IV.4 Mutational paths and bottleneck characterization
The mutational distance $M$ between ${\bf a}_{R}$ and ${\bf a}_{B}$ is defined as the number of differing sites. All $M!$ possible evolutionary paths between the reference variants are considered, and a value $E_{C}$ is determined such that at least one path of single mutations exists, maintaining $|E({\bf a})|>E_{C}$ at all intermediate steps. The distribution of $E_{C}$ and the likelihood of a single critical ‘jumper’ mutation define the presence and severity of the bottleneck.
IV.5 Data and Code availability
All simulations and analyses were performed using scripts available at Zenodo (DOI: 10.5281/zenodo.18131102). Data were downloaded from previously published sources Poelwijk et al. (2019); Buda et al. (2023). Specifically, raw sequencing data were retrieved from the Sequence Read Archive (SRA) under BioProject accession PRJNA560590.
@@ Acknowledgements.
We warmly thank Sid Nagel and Martin Weigt for providing very useful feedback and inspiration all along the development of this work, and Emily Hinds for very useful guidance in reproducing the data analysis from Ref. Poelwijk et al. (2019). This research has been supported by first FIS (Italian Science Fund) 2021 funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from MUR, Italian Ministry of University and Research and from the PRIN funding scheme (2022LMHTET - Complexity, disorder and fluctuations: spin glass physics and beyond) from MUR, Italian Ministry of University and Research.
Supplemental Material for
“Functional bottlenecks can emerge from non-epistatic underlying traits”
@@
The experimental measurements that were analyzed in the main text were obtained from the public repository associated with Ref. Poelwijk et al. (2019). These are raw sequencing data capturing the genotype variants present in the red and blue channel after two-color Fluorescence Activated Cell Sorting. Following the procedure detailed in Ref. Poelwijk et al. (2019) (see the original paper for details), sequences with low quality scores and noisy measurements were filtered out. The enrichment scores for each genotype a were computed counting the occurrence of each variant in the two channels normalized to its abundance in the input channel, i.e. before sorting. We further normalized the enrichment scores with respect to the ones of the reference variants, such that $F_{R}(\textbf{a}_{R})=F_{B}(\textbf{a}_{B})=1$, and afterwards removed global epistasis via a non-linear function
@@
as proposed by the authors of Ref. Poelwijk et al. (2019), separately for the blue (f=B{f}=B) and red (f=R{f}=R) fluorescence.
@@
Figure S1: (a) Scatter plot of the enrichment scores in the blue and the red channel, $\epsilon_{B}$ and $\epsilon_{R}$, not normalized with respect to the reference values. For the blue reference variant $\epsilon_{B}=1.58$, $\epsilon_{R}=0.09$. For the red one $\epsilon_{B}=0.17$, $\epsilon_{R}=0.73$. (b) Histogram of the values of $\widetilde{E}$ obtained for each of the $2^{13}$ variants. The red and blue lines correspond respectively to the reference values $\widetilde{E}^{\rm{ref}}_{R}$ and $\widetilde{E}^{\rm{ref}}_{B}$. (c) Topology of the space of paths obtained keeping only genotypes $\mathbf{a}$ with $|\widetilde{E}(\mathbf{a})|>\widetilde{E}_{C}$, where $\widetilde{E}_{C}=0.56$ (dashed lines) is the largest possible value such that the red and blue reference variants remain connected. The lower panel reports the values of $\widetilde{E}(\mathbf{a})$ for each functional genotype (blue dots for $\widetilde{E}>\widetilde{E}_{C}$ and red dots for $\widetilde{E}<-\widetilde{E}_{C}$) as a function of the number of mutations $m$ from the blue reference sequence, with the gray lines connecting pairs of genotypes that differ by a single mutation. The upper panel shows the resulting graph of connections.
@@
In the main text, and as discussed in Sec. S1.1, we normalized the enrichment scores to the reference values, a choice that was not made in the original Ref. Poelwijk et al. (2019). Using normalized enrichment scores, we find that no path of single mutations exists that connects the two reference variants while always maintaining $|E(\textbf{a})|\geq E^{\rm{ref}}_{\max}$, hence we need to consider $E_{\rm{th}}<E^{\rm{ref}}_{\max}$ in order to preserve the connection between the reference variants, as discussed in the main text. Instead, Ref. Poelwijk et al. (2019) shows a topology of the space of paths connecting genotypes with $|E(\textbf{a})|\geq E_{\rm{th}}$ and $E_{\rm{th}}=E^{\rm{ref}}_{\rm{min}}$. This discrepancy arises entirely from the choice of not normalizing the phenotypic trait.
[…]
@@
Compared to the normalized case shown in the main text, the resulting distribution of phenotypes, shown in Fig. S1b, is not symmetric around zero: the red reference sequence has $|\widetilde{E}^{\rm{ref}}_{R}|=0.56$ and the blue one is more than twice as large, with $\widetilde{E}^{\rm{ref}}_{B}=1.49$. This asymmetry arises entirely from the absence of normalization. When looking for the largest value of the functionality threshold $\widetilde{E}_{\rm{th}}$ such that at least one viable mutational path exists between the two reference variants, we find that $\widetilde{E}_{C}$ saturates to $\widetilde{E}^{\rm{ref}}_{\rm{min}}$ and does not produce a topology with a single jump. The topology is the same as the one shown in Ref. Poelwijk et al. (2019), and we consistently find that $\widetilde{E}_{C}$ = $\widetilde{E}^{\rm{ref}}_{\rm{min}}$ is the critical value above which any connection between the two reference sequences is lost. This is exactly what we would expect given the imbalance between the two reference phenotypes. However, given this imbalance, we believe that choosing a symmetric threshold, i.e. choosing the same $\widetilde{E}_{\rm{th}}$ for both functions, is not appropriate.
@@
We therefore believe that the choice of normalizing the enrichment scores to the reference variants provides a better view of the global fitness landscape, with the distribution of phenotypic values centered around zero. This choice results in a critical threshold $\widetilde{E}_{C}<\widetilde{E}_{\rm{ref}}$, as discussed in the main text.